OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. S(1/2, 1), T(1/2, 4) x = ?

OpenStudy (anonymous):

remind me what is standart form: Ax+By=C?

OpenStudy (anonymous):

Yes, or you could find the slope and use point slope form: \[y=m(x-x1) + y1\]

OpenStudy (anonymous):

where x1 and y1 could be 1/2 and 1

OpenStudy (anonymous):

ok, then first find the direction vector of the line: v=T-S=(1/2-1/2,4-1)=(0,3) then take perpedicular to it vector u=(-3,0) -3(x-1/2)+0(y-1)=0 -3x=-3/2 3x=3/2 this is the equation or in more simple form x=1/2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

yw